OpenStudy (anonymous):

Points U, V, and W are plotted on a number line. The coordinate of point V is –2. Points U and W are the same distance from point V. What could be the coordinates of U and W? A.U = –11, W = 7 B.U = –8, W = –6 C.U = –6, W = 6 D.U = 0, W = 2 free medal and fan

OpenStudy (anonymous):

those boxes are negative signs

OpenStudy (anonymous):

tbh idk this one ;-;

OpenStudy (anonymous):

tbh?

OpenStudy (anonymous):

to be honest

OpenStudy (anonymous):

just for trying

OpenStudy (anonymous):

ty I'll see if a friend of mine can help @undeadknight26

OpenStudy (anonymous):

ok tank you

OpenStudy (tkhunny):

Have you considered subtraction? A.U = –11, W = 7 (|-11-(-2)|) vs. |7-(-2)| -- Do they match B.U = –8, W = –6 (|-8-(-2)|) vs. |-6-(-2)| -- Do they match C.U = –6, W = 6 (|-6-(-2)|) vs. |6-(-2)| -- Do they match D.U = 0, W = 2 (|0-(-2)|) vs. |2-(-2)| -- Do they match